# Membership Inference Attacks and Privacy in Topic Modeling

### v1

@@ 3.5 Attack Evaluation Set-Up
We evaluate our attack against three data sets: TweetRumors, 20Newsgroup and NIPS [^1] The source for each dataset is included in the Availability section.. We apply standard text pre-processing procedures for each data set: removal of all non-alphabetic characters, tokenization, removal of stop words, removal of tokens longer than 15 characters, removal of tokens shorter than 3 characters, and lemmentization. [^2] Tokenization, stop words, and lemmentization implemented via python package nltk https://www.nltk.org Appendix C provides the basic data profile for each data set after pre-processing.
To initiate the attack, we randomly sample half of the data to learn and release $\Phi_{obs}$. Next, we train $N$ shadow topic models by repeatedly sampling half of the data set to simulate sampling from $\mathbb{D}$. In this scenario, the shadow model training data and the target model’s training data likely overlap. This is a strong assumption made to accommodate the smaller size of our data sets. As observed by [4], we do not expect attack performance to drop significantly using disjoint data sets.
We compare our LiRA against each of the attacks presented by [18]. To replicate their attacks, we randomly sample half of the data set to learn $\Phi_{obs}$. Using $\Phi_{obs}$, we estimate each documents’ topic-distribution $\hat{\theta}$ and compute the maximum posterior, standard deviation, and entropy on $\hat{\theta}$. We directly threshold the each of these statistics to evaluate membership for every document in the data set.
For each experiment, we learn $\Phi$ using scikit-learn’s implementation of LDA with default learning parameters. [^3] https://scikit-learn.org/ For TweetSet we set the number of topics $k=5$, 20Newsgroup we set $k=20$, and for NIPS we vary $k$. We learn $N=128$ shadow models, replicate each experiment 10 times and report our results across all iterations.
We interpret $\Lambda$ as a predicted membership score where a higher value indicates that the document is more likely to be a member of the training data. We empirically estimate the attack’s TPR at low FPRs and plot the attack’s ROC curve on log-scaled axes.
@@ 3.6 Attack Evaluation Results
First, we compare our online attack performance against the attacks in [18] at an FPR of 0.1% in Table 1. Figure 2 displays the ROC curves for the online and offline variant of our attacks. Table 1 and Figure 2 demonstrate that our LiRA outperforms each of attacks in [18].
Because the LiRA considers the likelihood-ratio of $d$, it accounts for per-example hardness and dominates the attack in Huang et al [18] at all FPR’s. As noted by [4], attacks that directly apply global thresholds do not consider document level differences on the learned model and fail to confidently identify members of the training data. Stronger attacks, like our LiRA, should be used to empirically evaluate the privacy associated with topic models.
@@
Data Set Our Attack Huang et al. [18]
@@
To understand how the number of topics in $\Phi$ influence attack performance, we vary the number of topics $k$ on NIPS and attack the resulting model. Figure 3 shows the ROC of the offline attack while varying $k$. As $k$ increases, we see that TPR increases at all FPRs. Furthermore, based on the Table 2, we note minor differences between online and offline attack performance.
Figure 3: Offline ROC Analysis While Varying $k$ on NIPS.
@@
The number of parameters in $\Phi$ grows linearly by the length of the vocabulary set as $k$ increases. With more topics and more parameters, document’s word co-occurrence structure is typically better represented in the document’s learned topic distribution. Consequently, increasing $k$ enhances the impact of including the document in the training data on the likelihood of a document resulting in better attack performance.
@@ 4.1 Differential Privacy
Definition 4.1 (Differential Privacy [11, 10])
Let $M:\mathcal{X}^{n}\rightarrow\mathcal{R}$ be a randomized algorithm. For any $\varepsilon\geq 0$ and $\delta\in[0,1]$, we say that $M$ is $(\varepsilon,\delta)$-differentially private if for all neighboring databases $x,x^{\prime}\in\mathcal{X}^{n}$ and every $S\subseteq\mathcal{R}$
@@
The notion of adjacency is critical. For any text data set, we say that two corpora $D$ and $D^{\prime}$ are author-level adjacency if they differ by one author’s documents. This notion enforces DPs promise of individual level privacy by considering adjacency at the user level.
@@
To achieve differential privacy we must add noise whose magnitude is scaled to the worst-case sensitivity of the target statistics from all adjacent data sets. Researchers have explored word-level adjacency to control sensitivity, but this approach offers weak privacy guarantees. For instance, most DP collapsed Gibbs sampling algorithms for learning LDA rely on word-level adjacency [33, 19, 18].
There are a variety of other algorithms for DP topic modeling: [22] propose DP stochastic variational inference to learn LDA, [9] use a privatized spectral algorithm to learn LDA, and [31] satisfy DP by adding noise directly to $\Phi$. These algorithms satisfy various notions of DP and use differing notions of adjacency. Table 5 in Appendix E summarizes each DP topic modeling implementation.
@@ 4.2 DP Vocabulary Selection
Releasing the topic-word distribution $\Phi$ without ensuring the privacy of the accompanying vocabulary set can lead to privacy violations. Topic models with comprehensive vocabulary sets tend to include many infrequently used words, and section 3.4 shows that documents with infrequently used words are more vulnerable to MIAs. Furthermore, even if the values in $\Phi$ are private, the overall release of the topic-word distribution can not satisfy DP because the vocabulary set accompanied by $\Phi$ is not private.
@@
DP vocabulary selection can be formalized as the DP Set-Union (DPSU) [13]. In the DPSU each author $i$ contributes a subset $W_{i}\subseteq U$ of terms in their documents, and our goal is to design a $(\varepsilon,\delta)$-DP algorithm to release a vocabulary set $S\subseteq\cup_{i}W_{i}$ such that the size of $S$ is as large as possible. Carvalho et al. propose a solution for DPSU which is particularly well fit for vocabulary selection because their method allows for one user to contribute the same term multiple times [7].
@@ 4.4 FDPTM Evaluation Set-Up
We empirically evaluate our FDPTM using the TweetRumors data set with $k=5$. We ensure author-level privacy via document-level adjacency under the assumption that each document has a unique author. For vocabulary selection we use the DPSU solution proposed in [7]. We use the DP LDA learning algorithm proposed by [34]. For DPSU, we fix the privacy parameter $\delta=10^{-5}$ and choose the cut-off value $\Gamma$ with the parameter $\alpha=3$, following [7].
To evaluate utility, we first vary the privacy loss parameter for DPSU and analyze utility when LDA is non-private. We study the interaction between DP vocabulary selection and DP LDA by fix the privacy loss parameter for one mechanism and vary the privacy loss parameter for the other mechanism. Specifically, we first fix $\varepsilon_{2}=3$ for DP LDA and vary $\varepsilon_{1}$ for DPSU. Then, we fix $\varepsilon_{1}=3$ and vary $\varepsilon_{2}$.
We evaluate the utility of the FDPTM using topic coherence [20], which can be a useful proxy for measuring the interpretability of topics. Topic coherence for a topic $t$ is defined as
@@
where $D(v)$ is the document frequency of word $v$ (i.e. the number of documents where word $v$ occurs), $D(v,v^{\prime})$ is the co-document frequency of words $v$ and $v^{\prime}$ (i.e. the number of documents where $v$ and $v^{\prime}$ both occur), and $V^{(t)}=\{v^{(t)}_{1},...,v^{(t)}_{M}\}$ is list of the $M$ most probable words in topic $t$. In our analysis, we select each topic’s top $M=10$ words and report the average across each topic.
@@ 4.5 FDPTM Evaluation Results
Figure 6 presents the ROC curves that explore the how the interaction between DP vocabulary selection and DP LDA affect attack performance. We gain some empirical privacy under FDPTM by decreasing LiRA performance at all FPRs. Generally, we see that attack performance decreases similarly as we decrease either privacy loss parameter. Furthermore, we gain empirical privacy by decreasing LiRA performance at all FPRs.
Figure 6: Attack ROCs while varying privacy loss parameters for DPSU and DP LDA. On the right-hand side, we fix $\varepsilon_{2}=5$ for DP LDA and vary $\varepsilon_{1}\in\{1,5,10\}$ for DPSU. On the left-hand side, we fix $\varepsilon_{2}$ and vary $\varepsilon_{1}$ at the same intervals. We provide the non-private baseline in red for reference.
@@
Our results indicate that dedicating most of the privacy budget to DP LDA, rather than DPSU, increases utility at the same privacy level. Intuitively, we can inject less noise into DP LDA, and more into the vocabulary selection algorithm ($\varepsilon_{1}<\varepsilon_{2}$) to increase model interpretability at the same global privacy loss ($\varepsilon_{1}+\varepsilon_{2}$) and similar empirical privacy against the LiRA. This approach allows us balance model privacy and utility in topic modeling.
@@ 5 Conclusions
To combat MIAs and memorization, we propose a modular framework for implementing better private topic models using differential privacy. Overall, the addition of DP vocabulary selection to the DP topic modeling work-flow is important for guaranteeing private, interpretable releases of $\Phi$. Not only does DP vocabulary selection allow for fully DP releases of $\Phi$, it also provides an effective defense against MIAs.
We highlight the greater need for continued development in the field of privacy-preserving ML. If simple probabilistic topics models for text memorize their training data, then it’s only inevitable that large language models and neural topic models memorize their training data. As ML models continue to become more sophisticated and widely used, privacy concerns become increasingly relevant.
@@
Additionally, our utility assessments for FDPTM are constrained by the limitations of automated topic evaluation methods. While different measures give us different insights into the topic model, the best evaluations are typically based on human inspection [17]. Other utility metrics that evaluate topic models with metrics that estimate model fit, like perplexity, tend to be negatively correlated with human-measured topic interpretability [8]. Therefore, we use topic coherence because it is associated with the human interpretability of topics.
Finally, our research focuses on cases where the topic modeling output is the entire topic-word distribution $\Phi$. In many practical use cases, researchers will release the top-$n$ words with or without word probabilities associated with each topic, word clouds of the top $n$ words in each topic, interactive visualizations such as LDAvis, or simply use topic modeling as dimensionality reduction [26]. Determining the affect that alternative releases have on attack performance by examining ways that clever adversaries can reconstruct $\Phi$ given each release mechanism is left for future research. We opt to attack $\Phi$ directly because it serves as the starting point for a variety of downstream tasks such as inferring unseen documents’ topic distributions for clustering or classification.
@@ Appendix A The Online LiRA
The LiRA on topic models follows directly from the LiRA on supervised ML models proposed by [4]. However, the attack on topic models differs because we use statistic $\zeta$ queried on $\Phi$. Algorithm 2 contains the online LiRA.
@@ Appendix B The Offline LiRA
The offline LiRA on topic models also follows directly from the offline LiRA on supervised ML models proposed by [4]. The offline variant benefits over the online variant because lines 3-9 can be executed regardless of the target document $d$, and lines 11-17 can be repeated for any $d$ after estimating $\textbf{T}_{out}$ once. Additionally, note that the test statistic $\Lambda$ changes in line 15. Algorithm 3 contains the offline LiRA.
@@
Appendix C Data Profile
Table 3: Data Set Profile After Pre-Processing
TweetRumors 5,698 9 5,942
NIPS 1,494 893 10,346
20Newsgroup 18,037 84 74,781
Appendix D Evaluating Candidate Query Statistics
@@
Additionally, we compare the proposed statistic to alternative statistics based on [18] and show that our proposed statistic outperforms the alternatives.
Huang et al.[18] propose three statistics: the entropy of $\hat{\theta}$, the standard deviation of $\hat{\theta}$, and the maximum value in $\hat{\theta}$ where $\hat{\theta}$ is a document’s estimated topic distribution [18]. Each statistic requires estimating the document’s topic distribution $\hat{\theta}$. They estimate $\hat{\theta}$ using sampling techniques, but for these experiments, we choose to estimate $\hat{\theta}$ by performing the optimization from Eq. 3 for increased efficiency and consistency with the proposed statistic in Eq. 3.
To tailor each statistic to fit within the LiRA framework, we make some minor modifications to the statistics proposed in [18]. Because entropy naturally decreases when the document is included in the training data, we use the negative entropy to satisfy the first requirement. Additionally, because the maximum value in $\hat{\theta}$ is bound between [0,1] and tends to concentrate toward 1 for outliers, we apply the logit transformation to promote adherence to the normality requirement, as illustrated in [4].
@@
We evaluate the normality requirement by performing a Shipiro-Wilk test for each $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$ with control for the False Discovery Rate (FDR) using the Benjamini-Hochberg Procedure [24, 1]. The null hypothesis is that the data is drawn from a normal distribution. Table 4 contains the the number of tests rejected with FDR control for each statistic. Our analysis indicates that the distributions of $\tilde{\textbf{T}}_{in/out}(d)$ for our statistic tend to exhibit normality. We can expect to encounter fewer rejections as we increase $N$ (i.e. increase the sample size of $\tilde{\textbf{T}}_{in/out}(d)$).
Table 4: Shapiro-Wilk Tests with FDR Control for 6000 Tests on Each Statistic

### v2

@@ 3.5 Attack Evaluation Set-Up
We evaluate our attack against three datasets: TweetRumors, 20Newsgroup and NIPS [^1] The evaluation code and source for each dataset is included in the Availability section (Availability).. We apply standard text pre-processing procedures for each dataset: removal of all non-alphabetic characters, tokenization, removal of stop words, removal of tokens longer than 15 characters, removal of tokens shorter than 3 characters, and lemmentization. [^2] Tokenization, stop words, and lemmatization implemented via python package nltk https://www.nltk.org Appendix D provides the basic data profile for each dataset after pre-processing.
To initiate the attack, we randomly sample half of the data to learn and release $\Phi_{obs}$. Next, we train $N$ shadow topic models by repeatedly sampling half of the data to simulate sampling from $\mathbb{D}$. Like in Carlini et al. (2022), this setup creates some overlap between shadow model training data and target model training data which presents a strong assumption made to accommodate the smaller size of our datasets. We do not expect performance to significantly decrease using disjoint datasets as observed by Carlini et al. (2022).
We compare our LiRA against each of the attacks presented by Huang et al. (2022). To replicate their attacks, we randomly sample half of the dataset to learn $\Phi_{obs}$. Using $\Phi_{obs}$, we estimate each documents’ topic-distribution $\hat{\theta}$ and compute the maximum posterior, standard deviation, and entropy on $\hat{\theta}$. We directly threshold the each of these statistics to evaluate membership for every document in the dataset.
For each experiment, we learn $\Phi$ using scikit-learn’s implementation of LDA with default learning parameters. [^3] https://scikit-learn.org/ For TweetSet and 20Newsgroup we set the number of topics $k$ to 5 and 20 respectively based on the data’s known distribution across topics: TweetSet contains tweets collected on 5 major news events and 20Newsgroup contains newsgroup documents across 20 different newsgroups. For NIPS, we vary $k$ for experimentation purposes. In practice, $k$ is known by the attacker who can observe the dimensions of $\Phi$. We learn $N=128$ shadow models, replicate each experiment 10 times and report our results across all iterations.
We interpret $\Lambda$ as a predicted membership score where a higher value indicates that the document is more likely to be a member of the training data. We empirically estimate our attacks’ performance by evaluating the TPR at all FPRs and plot the attack’s ROC curve on log-scaled axes.
@@ 3.6 Attack Evaluation Results
First, we compare our online attack performance against the attacks in Huang et al. (2022) at an FPR of 0.1% in Table 1. Figure 2 displays the ROC curves for the online and offline variant of our attacks. Table 1 and Figure 2 demonstrate that our LiRA outperforms each of attacks in Huang et al. (2022).
Because the LiRA considers the likelihood-ratio of $d$, it accounts for per-example hardness and dominates the attacks by Huang et al. (2022) at all FPR’s. As noted by Carlini et al. (2022), attacks that directly apply global thresholds do not consider document level differences on the learned model and fail to confidently identify members of the training data. Stronger attacks, like our LiRA, should be used to empirically evaluate the privacy associated with topic models.
@@
Dataset Our Attack Huang et al. Huang et al. (2022)
@@
To understand how the number of topics in $\Phi$ influence attack performance, we vary the number of topics $k$ on NIPS and attack the resulting model. Table 2 shows the the attack performance as we vary $k$. As $k$ increases, we see that TPR increases at an FPR of 0.1%. Additionally, we note minor differences between online and offline attack performance.
@@
The number of parameters in $\Phi$ grows linearly by the length of the vocabulary set as $k$ increases. With more topics and more parameters, document’s word co-occurrence structure is typically better represented in the document’s learned topic distribution. Consequently, increasing $k$ enhances the impact of including the document in the training data resulting in better attack performance.
@@ 4.1 Differential Privacy
Definition 4.1 (Differential Privacy (Dwork et al., 2006b; a))
Let $M:\mathcal{X}^{n}\rightarrow\mathcal{R}$ be a randomized algorithm. For any $\varepsilon\geq 0$ and $\delta\in[0,1]$, we say that $M$ is $(\varepsilon,\delta)$-differentially private if for all adjacent databases $x,x^{\prime}\in\mathcal{X}^{n}$ and every $S\subseteq\mathcal{R}$
@@
The notion of adjacency is critical. For any text dataset, we say that two corpora $D$ and $D^{\prime}$ are author-level adjacent if they differ by one author’s documents. This notion enforces DPs promise of individual level privacy by considering adjacency at the user level.
@@
To achieve differential privacy we must add noise whose magnitude is scaled to the worst-case sensitivity of the target statistics from all adjacent datasets. Researchers have explored word-level adjacency to control sensitivity, but this approach offers weak privacy guarantees. For instance, most DP collapsed Gibbs sampling algorithms for learning LDA rely on perturbing word-count statistics by adding noise scaled to word-level adjacency (Zhao et al., 2021; Huang & Chen, 2021; Huang et al., 2022).
There are a variety of other algorithms for DP topic modeling: Park et al. (2016) propose DP stochastic variational inference to learn LDA, Decarolis et al. (2020) use a privatized spectral algorithm to learn LDA, and Wang et al. (2022) satisfy DP by adding noise directly to $\Phi$. These algorithms satisfy various notions of DP and use differing notions of adjacency. Table 5 in Appendix E summarizes each DP topic modeling implementation.
@@ 4.2 DP Vocabulary Selection
Releasing the topic-word distribution $\Phi$ without ensuring the privacy of the accompanying vocabulary set can lead to privacy violations because the vocabulary set is derived directly from the data. Topic models with comprehensive vocabulary sets tend to include many infrequently used words from the corpus which makes them more vulnerable to MIAs (as shown in section 3.4). Furthermore, even if the values in $\Phi$ are private, the overall release of the topic-word distribution can not satisfy DP because the vocabulary set accompanied by $\Phi$ is not private.
@@
DP vocabulary selection can be formalized as the DP Set-Union (DPSU) (Gopi et al., 2020). If we assume each author $i$ contributes a subset $W_{i}\subseteq U$ of terms in their documents, DPSU seeks to design a $(\varepsilon,\delta)$-DP algorithm to release a vocabulary set $S\subseteq\cup_{i}W_{i}$ such that the size of $S$ is as large as possible. Researchers first approached this problem with the intent to release the approximate counts of as many items as possible in $\cup_{i}W_{i}$ (Korolova et al., 2009; Wilson et al., 2020). Their algorithms guarantee privacy by constructing a histogram of $\cup_{i}W_{i}$, adding noise to each word frequency in the histogram, and releasing all word frequencies that fall above a certain threshold. These implementations upper-bound the sensitivity based on the maximum the number of words contributed by all users ($\Delta_{0}=\max_{i}|W_{i}|$) to account for users contributing more than one word. However, most users contribute significantly less than $\Delta_{0}$ so this method wastes sensitivity. Gopi et al. (2020) propose algorithms that build weighted histograms with update policies that help control sensitivity. Intuitively, the update policy stops increasing the weights of words in the histogram after they reach some cutoff $\Gamma$ which is set using the parameter $\alpha$. Carvalho et al. (2022) propose an improved algorithm for DPSU which is well fit for vocabulary selection because it allows for one user to contribute the same term multiple times.
@@ 4.4 FDPTM Evaluation Set-Up
We empirically evaluate our FDPTM using the TweetRumors dataset with $k=5$. We ensure author-level privacy via document-level adjacency under the assumption that each document has a unique author. This assumption necessary for controlling sensitivity and is common to most DP NLP solutions. We accomplish this by using the DPSU solution proposed by Carvalho et al. (2022) for vocabulary selection and DP LDA learning algorithm by Zhu et al. (2016) because they satisfy the same notion of privacy. For DPSU, we fix the privacy parameter $\delta=10^{-5}$ and set the parameter $\alpha=3$ to control the cut-off value $\Gamma$ like in Carvalho et al. (2022).
To evaluate utility, we first vary the privacy loss parameter for DPSU and analyze utility when LDA is non-private. We study the interaction between DP vocabulary selection and DP LDA by fixing the privacy loss parameter for one mechanism and varying the privacy loss parameter for the other mechanism. Specifically, we first fix $\varepsilon_{2}=3$ for DP LDA and vary $\varepsilon_{1}$ for DPSU. Then, we fix $\varepsilon_{1}=3$ and vary $\varepsilon_{2}$. We vary our privacy loss parameters in intervals across common $\varepsilon$ choices for DP ML solutions. Typically, $\varepsilon>1$ provides very weak theoretical privacy guarantees, but tend to be effective for providing empirical privacy.
We evaluate the utility of the FDPTM using topic coherence (Mimno et al., 2011), which can be a useful proxy for measuring the interpretability of topics. Topic coherence for a topic $t$ is defined as
@@
where $D(v)$ is the document frequency of word $v$ (i.e. the number of documents where word $v$ occurs), $D(v,v^{\prime})$ is the co-document frequency of words $v$ and $v^{\prime}$ (i.e. the number of documents where $v$ and $v^{\prime}$ both occur), and $V^{(t)}=\{v^{(t)}_{1},...,v^{(t)}_{M}\}$ is list of the $M$ most probable words in topic $t$. A high topic coherence score suggests that the top $M$ terms have higher co-document frequencies which makes the topic easier to understand because the terms are more semantically similar or coherent. In our analysis, we select each topic’s top $M=10$ words and report the average across each topic.
@@ 4.5 FDPTM Evaluation Results
Figure 5 presents the ROC curves that explore the how the interaction between DP vocabulary selection and DP LDA affect attack performance. We gain some empirical privacy under FDPTM by decreasing LiRA performance at all FPRs. Generally, we see that attack performance decreases similarly as we decrease either privacy parameter. Furthermore, we gain empirical privacy by decreasing LiRA performance at all FPRs.
Figure 5: Attack ROCs while varying privacy loss parameters for DPSU and DP LDA. On the right-hand side, we fix $\varepsilon_{2}=5$ for DP LDA and vary $\varepsilon_{1}\in\{1,5,10\}$ for DPSU. On the left-hand side, we fix $\varepsilon_{1}$ and vary $\varepsilon_{2}$ at the same intervals. We provide the non-private baseline in red for reference.
@@
Our results indicate that dedicating most of the privacy budget to DP LDA, rather than DPSU, increases utility at the same privacy level. Intuitively, we can inject less noise into DP LDA, and more into the vocabulary selection algorithm ($\varepsilon_{1}<\varepsilon_{2}$) to increase model interpretability at the same global privacy loss ($\varepsilon_{1}+\varepsilon_{2}$) and similar empirical privacy against the LiRA. This approach allows us balance privacy and utility in topic modeling.
@@ 5 Conclusions
To combat MIAs and memorization, we propose a modular framework for implementing better private topic models using differential privacy. Overall, the addition of DP vocabulary selection to the DP topic modeling workflow is important for guaranteeing private, interpretable releases of $\Phi$. Not only does DP vocabulary selection allow for fully DP releases of $\Phi$, it also provides an effective defense against MIAs.
We highlight the greater need for continued development in the field of privacy-preserving ML. If simple probabilistic topics models for text memorize their training data, then it’s only inevitable that LLMs and neural topic models memorize their training data. As ML models continue to become more sophisticated and widely used, privacy concerns become increasingly relevant.
@@
Additionally, our utility assessments for FDPTM are constrained by the limitations of automated topic evaluation methods. While different measures give us different insights into the topic model, the best evaluations are typically based on human inspection (Hoyle et al., 2021). Other utility metrics that evaluate topic models with metrics that estimate model fit, like perplexity, tend to be negatively correlated with human-measured topic interpretability (Chang et al., 2009). Therefore, we use topic coherence because it is associated with the human interpretability of topics.
Finally, while our results are convincing, we conduct them with the same learning process ($f_{\mathcal{G}}$ based on LDA via Gibbs-sampling) and hyperparameters for the the shadow and target models with access to the entire topic-word distribution $\Phi$. In many practical use cases, researchers may not use LDA or release all of $\Phi$. A mismatch between shadow and topic models $f_{\mathcal{G}}$ may impact attack accuracy, but we suspect to a lesser extent than in neural models. Additionally, we believe that clever adversaries can reconstruct $\Phi$ given partial or alternative releases of $\Phi$. Determining the affect that different $f_{\mathcal{G}}$ and releases of $\Phi$ have on attack performance is left for future research.
@@ Appendix A The Online LiRA
The LiRA on topic models follows directly from the LiRA on supervised ML models proposed by Carlini et al. (2022). However, the attack on topic models differs because we use statistic $\zeta$ queried on $\Phi$. Algorithm 2 contains the online LiRA.
@@ Appendix B The Offline LiRA
The offline LiRA on topic models also follows directly from the offline LiRA on supervised ML models proposed by Carlini et al. (2022). The offline variant benefits over the online variant because lines 3-9 can be executed regardless of the target document $d$, and lines 11-17 can be repeated for any $d$ after estimating $\textbf{T}_{out}$ once. Additionally, note that the test statistic $\Lambda$ changes in line 15. Algorithm 3 contains the offline LiRA.
@@
Appendix C Evaluating Candidate Query Statistics
@@
Additionally, we compare the proposed statistic to alternative statistics based on the attacks in Huang et al. (2022) and show that our proposed statistic outperforms the alternatives.
Huang et al. (2022) propose three statistics: the entropy of $\hat{\theta}$, the standard deviation of $\hat{\theta}$, and the maximum value in $\hat{\theta}$ where $\hat{\theta}$ is a document’s estimated topic distribution Huang et al. (2022). Each statistic requires estimating the document’s topic distribution $\hat{\theta}$. They estimate $\hat{\theta}$ using sampling techniques, but for these experiments, we choose to estimate $\hat{\theta}$ by performing the optimization from Eq. 3 for increased efficiency and consistency with the proposed statistic in Eq. 3.
To tailor each statistic to fit within the LiRA framework, we make some minor modifications to the statistics proposed by Huang et al. (2022). Because entropy naturally decreases when the document is included in the training data, we use the negative entropy to satisfy the first requirement. Additionally, because the maximum value in $\hat{\theta}$ is bound between [0,1] and tends to concentrate toward 1 for outliers, we apply the logit transformation to promote adherence to the normality requirement, as illustrated by Carlini et al. (2022).
@@
We evaluate the normality requirement by performing a Shipiro-Wilk test for each $\tilde{\textbf{T}}_{in}(d)$ and $\tilde{\textbf{T}}_{out}(d)$ with control for the False Discovery Rate (FDR) using the Benjamini-Hochberg Procedure (Shapiro & Wilk, 1965; Benjamini & Hochberg, 1995). The null hypothesis is that the data is drawn from a normal distribution. Table 3 contains the the number of tests rejected with FDR control for each statistic at a significance level of .05. Our analysis indicates that the distributions of $\tilde{\textbf{T}}_{in/out}(d)$ for our statistic tend to be normal more often than the alternatives. We can expect to encounter fewer rejections as we increase $N$ (i.e. increase the sample size of $\tilde{\textbf{T}}_{in/out}(d)$).
[…]
@@
Appendix D Data Profile
Table 4: Dataset Profile After Pre-Processing
TweetRumors 5,698 9 5,942
NIPS 1,494 893 10,346
20Newsgroup 18,037 84 74,781